# Defining a Role-Centered Terminology for Physical Representations and Controls

### v1

@@ Abstract.
terminology, linguistics, blended words, physical computing, physical user interfaces, graspable user interfaces, tangible user interfaces, embodied interfaces, embedded interfaces, taxonomy, classification, role-centered viewpoint, what–how terminology, tangibility classes of applications
@@ 1. Introduction
Conceptualization often precedes a community’s focus on the same research path. Whereas precursor specimens—of what was later conceptualized as Tangible User Interfaces (TUIs)—arose in the mid-1970s and early 1980s Shaer and Hornecker, 2010: Ch. 2, Ullmer et al., 2022, physical computing came to the forefront of Human–Computer Interaction (HCI) researchers’ attention only fifteen years later (Wellner, 1991; Hinckley et al., 1994; Fitzmaurice et al., 1995), after the concept of ubiquitous computing was conceptualized in the late 1980s (Weiser, 1991, 1999). Several works initiated physical computing’s exploration, from “passive real-world props” (Hinckley et al., 1994), to “graspable user interfaces” (Fitzmaurice et al., 1995), to “manipulative user interfaces” (Harrison et al., 1998), to “embodied user interfaces” (Fishkin et al., 1999), which were the growing roots of the same tree (Fishkin, 2004). Likewise, the deeper exploration of the design and technology of tangible user interfaces began after they were conceptualized in 1997 (Ullmer and Ishii, 1997). Since then, while the number of specimens has grown yearly, taxonomies, paradigms, and frameworks have emerged Hoven and Eggen, 2004, Mazalek and Hoven, 2009, Shaer and Hornecker, 2010: Ch. 5, defining terms to describe the roles of physical artifacts (e.g., phicon (Ullmer and Ishii, 1997), token (Holmquist et al., 1999; Shaer et al., 2004; Ullmer et al., 2005), and tangible representation (Ullmer and Ishii, 2000)). However, no global terminology is shared in the literature, even if refining and conceptualizing tangible user interfaces’ interactive components better could help refine the design space, deepen specimens’ exploration, and broaden applications’ possibilities—indeed, sharing a common vocabulary and the same concepts also helps aggregating research efforts on a common path.
@@
Previous work provided isolated terms, common terms, and general terms, but none can distinguish all the entities involved in tangible user interfaces in a single namespace. On the one hand, isolated terms are limited in the scopes and paradigms that define them (e.g., Token And Constraints (Shaer et al., 2004)), and, therefore, remain separated in diverse namespaces. On the other hand, some terms are found across several research works, but with inconsistent meanings (e.g., “tokens” (Holmquist et al., 1999; Shaer et al., 2004; Ullmer et al., 2005)). As a consequence, many researches are built upon general terms (e.g., “tangibles,” “physical object,” and “tangible object” (Shaer and Hornecker, 2010)), thus escaping terminological ambiguity and describing all physical artifacts as a whole, without distinction. However, such general terms remain broad and undescriptive. For instance, the term “tangibles” is plural and cannot name merely a single artifact. The terms “physical object” and “tangible object” lend a new meaning to the term “object”, which is already overused in computer science (Shaer et al., 2004). Finally, “objects” are already “physical” and “tangible” in the usual sense.
@@
Furthermore, scopes other than pure tangible user interfaces resort to physical entities as representations of controls, but not especially for data representation (e.g., in Augmented Reality (Cordeil et al., 2020) and Virtual Reality (Mahieux et al., 2022) scopes). The lack of precise terms can lead to consider any user interface resorting to physical objects as a “tangible user interface,” even if these objects are only a subpart of a user interface that does not represent data with physical bodiement, thus leading to confusion and misunderstanding (e.g., before general frameworks, recognizing tangible user interfaces was mainly relying on “I know one when I see one” approaches (Fishkin, 2004)). Fishkin’s taxonomy overcame these considerations by “treating tangibility as a spectrum rather than a binary quantity” (i.e., “tangible” and “not tangible” user interfaces) and organizing “proof of concept” specimens over a twenty-item discretized space (Fishkin, 2004). However, this taxonomy, which has contributed to better delineating and understanding the field, does not address tangible user interfaces at the applicative level (i.e., user interfaces that gather and assemble components to meet the needs of final users). Indeed, this taxonomy is mainly focused on analyzing data representations (i.e., distance between data output and the inputs enabling data manipulation and edition, and the user actions’ degree of analogy to perceive and manipulate data) and describes user interfaces as wholes (i.e., without refining the components that compose applicative user interfaces). Except that such a global description requires determining what is “the” datum of the user interface, but this often becomes tricky in applicative user interfaces, which involve multiple tasks, outputs, or users. Although describing the morphology and meaning of tangible user interfaces globally enabled refining the concept of “tangibility” in physical user interfaces, developing applications of such user interfaces may require describing components separately and beyond data representations.
For instance, the Urp specimen (Urban Planning Workbench) melts, on the surface of a table, models of buildings with editing and navigation tools and with measurement objects, and all of these are augmented with video-projection to display information from the computational simulation.
For instance, the Urban Planning Workbench (Urp) (Ben-Joseph et al., 2001; Ishii et al., 2002; Underkoffler and Ishii, 1999) specimen enables the manipulation of building models on a surface table, along with editing and navigation tools, as well as measurement objects. All of these features are augmented with video projection to display information from the computational simulation.
@@
Describing this “workbench” application globally, as a whole, ignores all the subtleties of its functioning and, for example, the role played by intangible representations. As well, determining the distance between input and output fails when an applicative tangible user interface only conveys data.
@@ 2. Related Work
This section reviews related work on terms for physical representations and controls in tangible user interfaces, the creation of new words in the usual language through the word blending, and taxonomies and classifications of tangible user interfaces.
@@ 2.1. Terms for Physical Representations
The term “tangible” appeared early to describe a “tangible model” for 3D input in Computer-Aided Architectural Design (CAAD) (Aish and Noakes, 1984) and “tangible manipulation” in the Digital Desk (Wellner, 1991) (but the latter was later renamed “tactile interaction” (Wellner, 1993)). Later, the terms “graspable” and “tangible” were introduced to define Graspable User Interfaces (Fitzmaurice et al., 1995) and Tangible User Interfaces (Ullmer and Ishii, 1997). Finally, the term “intangible” appeared to name audio and video output modalities in the MCRit model [^1] MCRit: “Model-Control-Representation (intangible and tangible)” (Ullmer, 2002; Ullmer et al., 2005), first known as MCRpd: “Model-Control-Representation (physical and digital)” (Ullmer and Ishii, 2000). (Ullmer, 2002; Ullmer et al., 2005). Whereas the term “tangible” is used close to dictionary definitions, the term “intangible” differs from the usual sense: it simply means “not tangible.” Therefore, intangible representations are those generated by the computer; they may disappear when the computer is turned off (e.g., sound, video projection, and any audio and visual modalities or displays). This section summarizes the usual terms used in the literature to name the interactional components of tangible user interfaces, categorized into general, isolated, and shared terms, as well as terms employed in specific toolkits and design tools.
@@ 2.1.1. General Terms
Some terms are used in the literature to generically name the physical part of tangible user interfaces that embody digital information, including “physical artifacts” (Ullmer and Ishii, 2000), “physical representations” (Ullmer and Ishii, 2000), “tangible representations” (Ullmer, 2002; Ullmer et al., 2005), “tangibles” (Ullmer, 2002), and “tangible objects” (Shaer and Hornecker, 2010). Those terms are variations to name graspable and tangible interactional components broadly. However, the common use of the term “tangible” was initially defined for bodied data representations (i.e., not graspable anchors on tabletops). Finally, only one general term refers to neither graspable nor tangible components: “intangible representations” (Ullmer, 2002; Ullmer et al., 2005).
@@ 2.1.2. Isolated Terms
The term passive real-world prop (Hinckley et al., 1994) is used to name bodied representations of tools and data that map sensed user actions directly to control digital data displayed on a distant screen. Props can be symbolic (Ullmer and Ishii, 2000) (e.g., a sphere (Hinckley et al., 1994)) or iconic (Ullmer and Ishii, 2000) (e.g., a doll’s head (Hinckley et al., 1994)) data representations, combined with tools (e.g., a slicing plate (Hinckley et al., 1994)). The term “passive” refers to the system’s ability to track and reproduce movements made on the props. We can understand the choice of this term with the first loop of tangible interaction (Ishii, 2008): props provide a first passive feedback loop before computation’s feedback arises—the relative position of the props informs about cutting plane location without watching the screen (what is reinforced when using the doll’s head (Hinckley, 1996)).
The term physical handle (Fitzmaurice et al., 1995) (or phandle (Ullmer and Ishii, 1997)) refers to physical artifacts of general form (e.g., bricks and pucks) that can be attached to intangible shapes displayed on a surface. Once the physical handle acquires an intangible shape, motions are transmitted to control the digital object (e.g., position, orientation, and size).
The term phicon (Ullmer and Ishii, 1997) stands for “physical icons” of a specialized form that instantiates graphical user interface (GUI) components physically. Once put on a horizontal surface, phicons are linked to digital data (e.g., a building (Ullmer and Ishii, 1997)), and intangible data is updated to match phicons’ physical states (e.g., a displayed map is updated (Ullmer and Ishii, 1997)).
@@ 2.1.3. Shared Terms
Some terms do not name the artifacts by what they are but by their role and appear across several research works. For example, the terms “token,” “container,” and “tool” (Holmquist et al., 1999) define three roles in giving access to digital information. Tokens are specialized artifacts whose physical properties match the represented digital information (e.g., phicons (Ullmer and Ishii, 1997)). Containers are generic artifacts that can represent any type of information. Whereas tokens are continuously assigned to the same digital information, containers are associated with any digital information over time. Finally, tools are physical artifacts that have computational functions, enabling them to act on the displayed data (e.g., physical handles (Fitzmaurice et al., 1995)). Beyond naming physical artifacts, access points to digital information are also named: digital information is rendered by information faucets (e.g., display devices, speakers, and tactile devices). However, the “Token+Constraint” approach (Ullmer et al., 2005) suggests that all those artifacts are “tokens”: “containers” are “symbolic tokens with permanent binding,” “tokens” are “iconic tokens with dynamic binding,” and “tools” are “tokens bound to operations” (Ullmer et al., 2005).
@@
The “Token+Constraint” approach (Ullmer et al., 2005) reuses the term “token” in its usual sense, which is also its standard usage in computer science. The term “token” is then used in combination with the term “constraint,” where tokens are physical objects representing digital information, and constraints are physical regions bound to digital operations. Physical constraints structure the placement and combination of tokens, which systems can sense and interpret to trigger operations.
@@
The term token is also used in the usual sense in the “TAC Paradigm” (Shaer et al., 2004), based on the aforementioned “Token+Constraint” approach (Ullmer, 2002; Ullmer et al., 2005). However, the TAC Paradigm first names physical objects composing a tangible user interface under the term pyfo to move away from the various uses of the term “object.” Pyfo is a constructive and recursive term (not atomic): pyfos are made of pyfos (e.g., the six sides of a box are also pyfos (Shaer et al., 2004)). As well, pyfos are not limited to whole graspable physical objects: pyfos are also intangible physical objects (e.g., intangible constraints drawn on a surface, such as Senseboard’s displayed grid, are pyfos (Jacob et al., 2002)). Thus, tokens are graspable pyfos representing digital information or computational functions; constraints are graspable or intangible pyfos that structure the behavior of tokens. Beyond physical representations, TAC Paradigm’s terminology also includes terms for digital elements, such as the term variable, which names digital information or computational functions. Whereas some variables can be bound to tokens, semantic variables cannot. Finally, systems behave according to a series of rules: association triplets between a token, a variable, and some constraints. These rules are referred to as TAC (i.e., Token And Constraints).
@@ 2.1.4. Terms in Toolkits and Design Tools
The term phidgets (Greenberg and Fitchett, 2001) stands for “physical widgets”: packaged, low-level devices and software architecture inspired by GUI toolkits to ease the building of physical user interfaces. The Phidgets toolkit provides physical primitives for controllers, sensors, and mechanical parts that are commercially available [^2] Phidgets’ website: https://www.phidgets.com/, last accessed 2022-06-09..
ASUR++ (Dubois et al., 2002, 2003) is a notation for describing the physical and digital entities of mobile mixed systems. This notation extends the ASUR notation for mixed reality systems and is usable for tangible user interfaces (Dubois et al., 2003). The notation defines two terms for physical objects, by distinction of two digital roles: tools that support a task (Rtool) and targeted objects that are modified by a task (Robject).
The ROSS API (Wu et al., 2012) defines four nested levels to describe physical representations and controls. A strength of this low-level framework is that it goes beyond the description of physical objects (RObject) by denoting active spaces (RSpace) and surfaces (RSurface) that can sense interaction. However, even if input controls—such as buttons, switches, and sliders—are described (RControl), physical objects’ roles—such as tools and data representations—are not describable by the API.
@@ 2.1.5. Existent Terms’ Limitations
The previous terms are built from two main strategies: drawing a parallel with GUI elements (e.g., “icon,” “widget,” “handle,” and “container”) or naming their physical roles (e.g., “prop,” “token,” and “constraint”). At certain points, these approaches enable the description of low embodiment (e.g., “phandles”) and higher embodiment (e.g., “phicon”). However, these approaches fail to establish a global terminology that can describe all the components residing within tangible user interfaces. For example, “tokens” can equally name artifacts representing data or functions. Moreover, tangible user interfaces often comprise representations that are not “purely tangible” (i.e., not representing data), but no terms distinguish a pure tangible user interface from other related physical user interfaces. Thereby, the lack of descriptive and discriminative power leads to the wide use of general terms to name any artifact indistinctly.
@@
As a general term, this article considers that “physical representations” comprise tangible, graspable, and intangible representations (as pyfos that also consider intangible representations as being physical)—the difference between “tangible” and “graspable” representations residing in the degree of embodiment: high and low, respectively. High embodiment occurs when some of the digital entities’ representational characteristics are embedded in their representations (similarly to the specialized form artifact definition (Fitzmaurice and Buxton, 1997)) and have iconic meaning (Ullmer and Ishii, 2000). Low embodiment occurs when no digital entities’ representational characteristics are embodied by their representations (similarly to the generic form artifact’s definition (Fitzmaurice and Buxton, 1997)) and that have symbolic meaning (Ullmer and Ishii, 2000). This article also considers that the term “artifacts” refers to bodied objects, regardless of whether they are graspable or tangible. This article introduces a new holistic terminology that refines the descriptive level of the aforementioned physical representations by employing word-blending formation processes.

### v3

@@ Abstract.
terminology, linguistics, blended words, physical computing, physical user interfaces, graspable user interfaces, tangible user interfaces, embodied interfaces, taxonomy, classification
@@ 1. Introduction
Conceptualization often precedes a community’s focus on the same research path. Whereas precursor specimens—of what was later conceptualized as Tangible User Interfaces (TUIs)—arose in the mid-1970s and early 1980s shaer2010past: Ch. 2, ullmer2022weaving, physical computing came to the forefront of Human–Computer Interaction (HCI) researchers’ attention only fifteen years later (wellner1991digitaldesk; hinckley1994passiveprops; fitzmaurice1995bricks), after the concept of ubiquitous computing was conceptualized in the late 1980s (weiser1991century; weiser1999century). Several works initiated physical computing’s exploration, from “passive real-world props” (hinckley1994passiveprops), to “graspable user interfaces” (fitzmaurice1995bricks), to “manipulative user interfaces” (harrison1998manipulative), to “embodied user interfaces” (fishkin1999embodied), which were the growing roots of the same tree (fishkin2004taxonomy). Likewise, the deeper exploration of the design and technology of tangible user interfaces began after they were conceptualized in 1997 (ullmer1997metadesk). Since then, while the number of specimens has grown yearly, taxonomies, paradigms, and frameworks have emerged hoven2004tangible, mazalek2009framing, shaer2010past: Ch. 5, defining terms to describe the roles of physical artifacts (e.g., phicon (ullmer1997metadesk), token (holmquist1999token; shaer2004tac; ullmer2005token), and tangible representation (ullmer2000emerging)). However, no global terminology is shared in the literature, even if refining and conceptualizing tangible user interfaces’ interactive components better could help refine the design space, deepen specimens’ exploration, and broaden applications’ possibilities—indeed, sharing a common vocabulary and the same concepts also helps aggregating research efforts on a common path.
@@
Previous work provided isolated terms, common terms, and general terms, but none can distinguish all the entities involved in tangible user interfaces in a single namespace. On the one hand, isolated terms are limited in the scopes and paradigms that define them (e.g., Token And Constraints (shaer2004tac)), and, therefore, remain separated in diverse namespaces. On the other hand, some terms are found across several research works, but with inconsistent meanings (e.g., “tokens” (holmquist1999token; shaer2004tac; ullmer2005token)). As a consequence, many researches are built upon general terms (e.g., “tangibles,” “physical object,” and “tangible object” (shaer2010past)), thus escaping terminological ambiguity and describing all physical artifacts as a whole set, without distinction. However, such general terms remain broad and undescriptive. For instance, the term “tangibles” is plural and cannot name merely a single artifact. The terms “physical object” and “tangible object” lend a new meaning to the term “object”, which is already overused in computer science (shaer2004tac). Finally, “objects” are already “physical” and “tangible” in the usual sense.
@@
Furthermore, scopes other than pure tangible user interfaces resort to physical entities as representations of controls, but not especially for data representation (e.g., in Augmented Reality (cordeil2020axes) and Virtual Reality (mahieux2022sablier) scopes). The lack of precise terms can lead to consider any user interface resorting to physical objects as a “tangible user interface,” even if these objects are only a subpart of a user interface that does not represent data with physical bodiement, thus leading to confusion and misunderstanding (e.g., before general frameworks, recognizing tangible user interfaces was mainly relying on “I know one when I see one” approaches (fishkin2004taxonomy)). Fishkin’s taxonomy overcame these considerations by “treating tangibility as a spectrum rather than a binary quantity” (i.e., “tangible” and “not tangible” user interfaces) and organizing “proof of concept” specimens over a twenty-item discretized space (fishkin2004taxonomy). However, this taxonomy, which has contributed to better delineating and understanding the field, does not address tangible user interfaces at the applicative level (i.e., user interfaces that gather and assemble components to meet the needs of final users). Indeed, this taxonomy is mainly focused on analyzing data representations (i.e., distance between data output and the inputs enabling data manipulation and edition, and the user actions’ degree of analogy to perceive and manipulate data) and describes user interfaces as whole things (i.e., without refining the components that compose applicative user interfaces). Except that such a global description requires determining what is “the” datum of the user interface, but this often becomes tricky in applicative user interfaces, which often involve multiple tasks, outputs, or users. Although describing the morphology and meaning of tangible user interfaces globally enabled refining the concept of “tangibility” in physical user interfaces, developing applications of such user interfaces may require describing components separately and beyond data representations.
@@
For instance, the Urp specimen (Urban Planning Workbench) (benjoseph2001urp; ishii2002urp; underkoffler1999urp) melts, on the surface of a table, models of buildings with editing and navigation tools and with measurement objects, and all of these are augmented with video-projection to display information from the computational simulation. Describing this “workbench” application globally, as a whole thing, ignores all the subtleties of its functioning and, for example, the role played by intangible representations. As well, determining the distance between input and output fails when an applicative tangible user interface only conveys data (i.e., no input).
@@ 2. Related Work
This section reviews the related work on terms for physical representations and controls in tangible user interfaces, the creation of new words in the usual language through the word blending, and taxonomies and classifications of tangible user interfaces.
@@ 2.1. Terms for Physical Representations
The term “tangible” appeared early to describe a “tangible model” for 3D input in Computer-Aided Architectural Design (CAAD) (aish1984caad) and “tangible manipulation” in the Digital Desk (wellner1991digitaldesk) (but the latter was later renamed “tactile interaction” (wellner1993digitaldesk)). Later, the terms “graspable” and “tangible” were introduced to define Graspable User Interfaces (fitzmaurice1995bricks) and Tangible User Interfaces (ullmer1997metadesk), respectively. Finally, the term “intangible” appeared to name audio and video output modalities in the MCRit model [^1] MCRit: “Model-Control-Representation (intangible and tangible)” (ullmer2002phdthesis; ullmer2005token), first known as MCRpd: “Model-Control-Representation (physical and digital)” (ullmer2000emerging). (ullmer2002phdthesis; ullmer2005token). Whereas the term “tangible” is used close to dictionary definitions, the term “intangible” differs from the usual sense: it simply means “not tangible.” Therefore, intangible representations are those generated by the computer; they may disappear when the computer is turned off (e.g., sound, video projection, and any audio and visual modalities or displays).
This section summarizes the usual terms used in the literature to name the interactional components of tangible user interfaces, categorized into general, isolated, and shared terms, as well as terms employed in specific toolkits and design tools.
@@ 2.1.1. General Terms
Some terms are used in the literature to generically name the physical part of tangible user interfaces that embody digital information, including “physical artifacts” (ullmer2000emerging), “physical representations” (ullmer2000emerging), “tangible representations” (ullmer2002phdthesis; ullmer2005token), “tangibles” (ullmer2002phdthesis), and “tangible objects” (shaer2010past). Those terms are variations to name graspable and tangible interactional components broadly. However, the common use of the term “tangible” was initially defined for bodied data representations (i.e., not for graspable anchors on tabletops). Finally, only one general term refers to neither graspable nor tangible components: “intangible representations” (ullmer2002phdthesis; ullmer2005token).
@@ 2.1.2. Isolated Terms
The term passive real-world prop (hinckley1994passiveprops) is used to name bodied representations of tools and data that map sensed user actions directly to control digital data displayed on a distant screen. Props can be symbolic (ullmer2000emerging) (e.g., a sphere (hinckley1994passiveprops)) or iconic (ullmer2000emerging) (e.g., a doll’s head (hinckley1994passiveprops)) data representations, combined with tools (e.g., a slicing plate (hinckley1994passiveprops)). The term “passive” refers to the system’s ability to track and reproduce movements made on the props. We can understand the choice of this term with the first loop of tangible interaction (ishii2008beyond): props provide a first passive feedback loop before computation’s feedback arises—the relative position of the props informs about cutting plane location without watching the screen (what is reinforced when using the doll’s head (hinckley1996thesis)).
The term physical handle (fitzmaurice1995bricks) (or phandle (ullmer1997metadesk)) refers to physical artifacts of general form (e.g., bricks and pucks) that can be attached to intangible shapes displayed on a surface. Once the physical handle acquires an intangible shape, motions are transmitted to control the digital object (e.g., position, orientation, and size).
The term phicon (ullmer1997metadesk) stands for “physical icons” of a specialized form that instantiates graphical user interface (GUI) components physically. Once put on a horizontal surface, phicons are linked to digital data (e.g., a building (ullmer1997metadesk)), and intangible data is updated to match phicons’ physical states (e.g., a displayed map is updated (ullmer1997metadesk)).
@@ 2.1.3. Shared Terms
Some terms do not name the artifacts by what they are but by their role and appear across several research works. For example, the terms “token,” “container,” and “tool” (holmquist1999token) define three roles in giving access to digital information. Tokens are specialized artifacts whose physical properties match the represented digital information (e.g., phicons (ullmer1997metadesk)). Containers are generic artifacts that can represent any type of information. Whereas tokens are continuously assigned to the same digital information, containers are associated with any digital information over time. Finally, tools are physical artifacts that have computational functions, enabling them to act on the displayed data (e.g., physical handles (fitzmaurice1995bricks)). Beyond naming physical artifacts, access points to digital information are also named: digital information is rendered by information faucets (e.g., display devices, speakers, and tactile devices). However, the “Token+Constraint” approach (ullmer2005token) suggests that all those artifacts are “tokens”: “containers” are “symbolic tokens with permanent binding,” “tokens” are “iconic tokens with dynamic binding,” and “tools” are “tokens bound to operations” (ullmer2005token).
@@
The “Token+Constraint” approach (ullmer2005token) reuses the term “token” in its usual sense, which is also its standard usage in computer science. The term “token” is then used in combination with the term “constraint,” where tokens are physical objects representing digital information, and constraints are physical regions bound to digital operations. Physical constraints structure the placement and combination of tokens, which systems can sense and interpret to trigger operations.
@@
The term token is also used in the usual sense in the “TAC Paradigm” (shaer2004tac), based on the aforementioned “Token+Constraint” approach (ullmer2002phdthesis; ullmer2005token). However, the TAC Paradigm first names physical objects composing a tangible user interface under the term pyfo to move away from the various uses of the term “object.” Pyfo is a constructive and recursive term (not atomic): pyfos are made of pyfos (e.g., the six sides of a box are also pyfos (shaer2004tac)). As well, pyfos are not limited to whole graspable physical objects: pyfos are also intangible physical objects (e.g., intangible constraints drawn on a surface, such as Senseboard’s displayed grid, are pyfos (jacob2002senseboard)). Thus, tokens are graspable pyfos representing digital information or computational functions; constraints are graspable or intangible pyfos that structure the behavior of tokens. Beyond physical representations, TAC Paradigm’s terminology also includes terms for digital elements, such as the term variable, which names digital information or computational functions. Whereas some variables can be bound to tokens, semantic variables cannot. Finally, systems behave according to a series of rules: association triplets between a token, a variable, and some constraints. These rules are referred to as TAC (i.e., Token And Constraints).
@@ 2.1.4. Terms in Toolkits and Design Tools
The term phidgets (greenberg2001phidgets) stands for “physical widgets”: packaged, low-level devices and software architecture inspired by GUI toolkits to ease the building of physical user interfaces. The Phidgets toolkit provides physical primitives for controllers, sensors, and mechanical parts that are commercially available [^2] Phidgets’ website: https://www.phidgets.com/, last accessed 2022-06-09..
ASUR++ (dubois2002asur++; dubois2003asur++) is a notation for describing the physical and digital entities of mobile mixed systems. This notation extends the ASUR notation for mixed reality systems and is usable for tangible user interfaces (dubois2003asur++). The notation defines two terms for physical objects, by distinction of two digital roles: tools that support a task (Rtool) and targeted objects that are modified by a task (Robject).
The ROSS API (wu2012api) defines four nested levels to describe physical representations and controls. A strength of this low-level framework is that it goes beyond the description of physical objects (RObject) by denoting active spaces (RSpace) and surfaces (RSurface) that can sense interaction. However, even if input controls—such as buttons, switches, and sliders—are described (RControl), physical objects’ roles—such as tools and data representations—are not describable by the API.
@@ 2.1.5. Existent Terms’ Limitations
The previous terms are built from two main strategies: drawing a parallel with GUI elements (e.g., “icon,” “widget,” “handle,” and “container”) or naming their physical roles (e.g., “prop,” “token,” and “constraint”). At certain points, these approaches enable the description of low embodiment (e.g., “phandles”) and higher embodiment (e.g., “phicon”). However, these approaches fail to establish a global terminology that can describe all the components residing within tangible user interfaces. For example, “tokens” can equally name artifacts representing data or functions. Moreover, tangible user interfaces often comprise representations that are not “purely tangible” (i.e., not representing data and not with high embodiment), but no terms distinguish a “pure” tangible user interface from other related physical user interfaces. Thereby, the lack of descriptive and discriminative power leads to the wide use of general terms to name any artifact indistinctly.
@@
As a general term, this article considers that “physical representations” comprise tangible, graspable, and intangible representations (as pyfos that also consider intangible representations as being physical)—the difference between “tangible” and “graspable” representations residing in the degree of embodiment: high and low, respectively. High embodiment occurs when some of the digital entities’ representational characteristics are embedded in their representations (similarly to the specialized form artifact definition (fitzmaurice1997empirical)) and have iconic meaning (ullmer2000emerging). Low embodiment occurs when no digital entities’ representational characteristics are embodied by their representations (similarly to the generic form artifact’s definition (fitzmaurice1997empirical)) and that have symbolic meaning (ullmer2000emerging). This article also considers that the term “artifacts” refers to bodied objects, regardless of whether they are graspable or tangible. Finally, this article introduces a new holistic terminology that refines the descriptive level of the aforementioned physical representations by employing word-blending formation processes.